Assume the following, m1 = .15, i1 = .15, and c1 = .8. The increase in government expenditure of 4 would expect to increase national income by:

a) 20

b) 11.44

c) 16

d) 3.2
Answer #1

Multiplier = 1/(1-.8+.15-.15) = 5

Increase in national income = 5*4= 20

a) 20 - is correct